# BEVPR: BEV-Enhanced Visual Place Recognition with Structural Cues

### v1

@@ III-F Network Training
IV EXPERIMENTS
IV-A Dataset
Partitioning of Positive and Negative Samples. Given some characteristics that the spatial positions of consecutive frames are close and the camera angle changes are limited, while appearance such as illumination undergoes minimal variations, treating consecutive frames as positive samples proves excessively facile and lacks practical significance. Consequently, images within the same scene are not regarded as positive samples to one another. The selection of negative samples is straightforward: for each query sample, samples other than its positive samples and consecutive frames are defined as negative samples.
The exclusive reliance on GPS for distinguishing between positive and negative samples presents challenges. Images captured in close geographical location may fail to depict identical scenes when oriented differently, leading to pseudo positive samples. And variations in camera position while photographing the same site may introduce the possibility of pseudo negative samples. Due to the limitations in reflecting place position with camera coordinates, we employ image coordinates rather than camera coordinates for determining positive samples corresponding to a query sample.
In this work, the position of an image is defined as in front of the camera. In order to obtain more reasonable initial positive and negative samples, each image is represented by a binary tuple , where, represents the directional vector of the camera-captured image, denoted as
where represents the position of the camera.
Then, we apply the Euclidean distance metric to calculate the distance between two images and those with a distance less than the threshold from other images are regarded as candidate positive samples, followed by calculating the angle between the direction vectors of query and positive to determine the final positive samples, i.e.,
where and represent the direction vectors of the query image and the candidate positive images respectively.
Furthermore, in order to more clearly analyze the advantages and disadvantages of different modalities, we classify sample types based on the appearance differences between query and positive images, and define it as follows (refer to Table II for more details):
• Hard Sample: significant illumination difference. For instance, the query image is taken during daylight, while all positive images are captured at nighttime.
@@ IV-A Dataset
Partition of Training and Test Sets. Considering that NuScenes only involves four regions, each with data collected in relatively concentrated periods, we opt for a holistic partitioning approach for data across all areas. In VPR, the research on recalling from day to night, as well as from night to rain presents more challenges, but holds greater significance. Therefore, We aim to achieve a relatively uniform distribution of such data in both the training and testing sets. This data distribution design facilitates a more comprehensive evaluation of the model’s adaptability to complex environmental changes, enhancing the model’s generalization performance. Note that there is no any instance and scene overlap between the training set and the test set.
We first establish an undirected graph with scenes as nodes and similarity between scenes as edge weights, where the scene similarity is defined as the proportion of each query image in one scene that has a corresponding positive image appearing in another scene. Specifically, If all query images have positive images in the same scene, the scene similarity is set to 1. After processing all the scenes, there are 32 isolated scenes without connecting to any other scenes, making them unusable for VPR but usable for BEV generation. And the graph composed of other scenes involves 47 connected components. Finally, we leverage the official API of NuScenes dataset to acquire scene labels, such as "night" or "day", and based on which, the data is manually balanced, generating the partitioning results in Table I. Further details are presented in https://github.com/FudongGe/BEV2PR.
@@ IV-B Implementation Details
Architecture. We implement BEVPR in PyTorch framework and only reproduce open-source models for a fair comparison, including SPoC [25], NetVLAD [12], GeM [15], Conv-AP [18], EigenPlaces [26], MixVPR [27]. For all techniques involving BEV generation and VPR, we use EfficientNet-B0 as the image backbone cropped at the last convolutional layer as recommended by their authors. Considering the information advantages of different scales, we adopt a method of upsampling the output of the block and concatenating it with the output of the block as final features. In addition, the sub-backbone in visual stream is cropped from the block of EfficientNet-B0. The camera-to-BEV model outputs a BEV map at a resolution with three classes, including lane boundary (Bound.), lane divider (Div.), and pedestrian crossing (P.C.), i.e., . Each input front-view image is resized from to in all experiments for all models.
Training. During Stage I, we train the camera-to-BEV network using EfficientNet-B0 pre-trained on ImageNet for 40 epochs on four NVIDIA Tesla V100 GPUs. The Adam optimizer to update the weights of the network has a weight decay of with a learning rate of .
During Stage II, for maximum fairness in VPR, the final descriptor dimension of all methods is set to . We train the whole network using Adam optimizer having a weight decay of with a learning rate of for epochs. For each mini-batch, the number of positive samples , negative samples = 6 and are set for the triplet loss. And the constants in the multi-head loss function are set to , , . In addition, we propose a newly online adaptive hard mining strategy to avoid model collapse caused by unstable gradients when the hard ratio is too high, as well as the neglect of samples rich in information when the hard ratio is too low. To specify, the hard ratio is dynamically adjusted based on loss changes.
Evaluation. The recall at Top-K (K = 1, 5, 10) is adopted as the evaluation metric, based on which, we evaluate our framework on the overall dataset as well as three subsets with different levels of recall difficulty, i.e., R@K (overall), R@K (easy), R@K (hard), R@K (semi-hard).
@@
| Improvements | - | - | - | +1.60 | +0.63 | +0.69 | +0.91 | +0.13 | +0.02 | +10.03 | +13.18 | +16.05 | +3.14 | -0.63 | -0.42 |
| # | Arch | BEV |
R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | ||
| 1 |
w/o | unshared | 88.93 | 93.87 | 94.80 | 94.48 | 98.62 | 99.16 | 4.58 | 8.88 | 12.03 | 84.83 | 94.98 | 97.49 | ||
| 2 | Single-input | w/o | shared | 88.75 | 93.73 | 94.90 | 93.70 | 97.79 | 98.50 | 10.89 | 19.48 | 26.93 | 86.51 | 95.61 | 97.28 | ||
| Improvements | - | - | -0.18 | -0.15 | +0.10 | -0.78 | -0.84 | -0.66 | +6.30 | +10.60 | +14.90 | +1.67 | +0.63 | -0.21 | |||
| 3 | Dual-input | w | unshared | 90.59 | 94.03 | 95.45 | 95.24 | 98.21 | 99.14 | 15.19 | 22.64 | 29.23 | 89.12 | 94.04 | 96.55 | ||
| Improvements | - | - | +1.66 | +0.16 | +0.65 | +0.76 | -0.41 | -0.02 | +10.61 | +13.76 | +17.20 | +4.29 | -0.94 | -0.94 | |||
| 4 | Single-input | w | shared | 91.40 | 95.53 | 96.85 | 95.67 | 98.46 | 99.03 | 22.64 | 42.41 | 54.44 | 89.85 | 96.65 | 98.74 | ||
| Improvements | - | - | +2.47 | +1.65 | +2.05 | +1.19 | -0.17 | -0.13 | +18.05 | +33.52 | +42.41 | +5.02 | +1.67 | +1.26 |
@@ IV-C Evaluation for Visual Place Recognition
We apply six baseline models to our framework and evaluate the VPR performance based on monocular images. Table III shows the recall of methods on the VPR-NuScenes dataset. It can be seen that all methods incorporating structural streams have shown improvements in almost all metrics, with the absolute increments on R@1 across the overall dataset being 5.60%, 2.59%, 4.19%, 2.47%, 2.80%, 1.60%, and the ones on R@5 being 2.99%, 1.80%, 2.62%, 1.65%, 0.88%, 0.63%, respectively. We observe that the dual-stream models are relatively mediocre on the easy subset. Such results are due to the fact that models can achieve considerable performance using only appearance information under similar environments of illumination and weather, but also prove that datasets do not limit BEVPR: on datasets with significant appearance differences, it can effectively improve performance (refer to next two paragraphs); on datasets with indistinct appearance differences, BEVPR maintains RGB performance without hurting it.
@@
Illumination Change. Compared to the easy set, the performance improvement on the hard set is more significant, with a maximum increase of 18.05% on R@1. This can be attributed to the fact that in scenes with significant illumination changes, the semantic and texture information becomes completely ineffective and structural features in the BEV frame are relatively stable. As shown in the second row of Fig. 3, it’ s difficult for human eyes to distinguish the complete scene, while just like in the nighttime scenes of autonomous driving in the real world, grounds are particularly clear compared to other areas, especially lane boundary and lane divider. Fig. 4 presents the RGB and BEV feature map during the day and night, respectively. It is evident that under such conditions, the effectiveness of structural information in the BEV frame is far greater than that of appearance information in the RGB frame.
| 1 | - | - | - | 88.93 | 93.87 | 94.80 | 94.48 | 98.62 | 99.16 | 4.58 | 8.88 | 12.03 | 84.83 | 94.98 | 97.49 |
| 2 | ✓ | - | - | 91.25 | 94.90 | 96.05 | 96.09 | 98.83 | 99.41 | 14.33 | 25.50 | 32.95 | 89.12 | 95.71 | 98.12 |
| 3 | - | ✓ | - | 89.96 | 94.21 | 95.56 | 95.19 | 98.36 | 99.08 | 9.46 | 22.35 | 30.66 | 86.72 | 94.56 | 97.28 |
| 4 | - | - | ✓ | 90.11 | 94.90 | 96.20 | 94.94 | 98.84 | 99.46 | 15.76 | 24.93 | 34.96 | 87.13 | 95.82 | 98.22 |
| 5 | ✓ | ✓ | - | 91.02 | 95.05 | 96.30 | 96.09 | 98.59 | 99.30 | 14.33 | 30.09 | 39.54 | 87.34 | 96.65 | 98.33 |
| 6 | ✓ | - | ✓ | 91.11 | 94.65 | 95.82 | 95.92 | 98.59 | 99.20 | 17.48 | 26.07 | 33.24 | 87.97 | 95.08 | 97.59 |
| 7 | - | ✓ | ✓ | 90.77 | 94.69 | 95.68 | 95.79 | 98.84 | 99.48 | 9.74 | 17.19 | 23.21 | 89.02 | 97.07 | 98.43 |
| 8 | ✓ | ✓ | ✓ | 91.40 | 95.53 | 96.85 | 95.67 | 98.46 | 99.03 | 22.64 | 42.41 | 54.44 | 89.85 | 96.65 | 98.74 |
@@
Image Clarity Change. The methods with added structural cues can also yield considerable performance enhancements in scenes where there is a substantial change in image clarity, with an average improvement of 5.62% on R@1 across different models. As shown in the third row of Fig. 3, compared with positive images, the query image is more blurry, but the overall image content remains basically unchanged. The semantic and texture information in certain areas is relatively complete and the structural information can further assist in generating more comprehensive features.
These results demonstrate the importance of BEV representation in compensating for deficiencies in VPR.
@@ IV-D Ablation Study
We perform all ablation experiments with Conv-AP as the baseline model.
Shared Bottom Backbone. To demonstrate the effectiveness of the shared bottom backbone, we intentionally degenerate our framework into two versions, as depicted in Table IV, one (#2) is single-input with a bottom backbone pre-trained through BEV generation but without structural stream, the other (#3) is dual-input with a bottom backbone not pre-trained via BEV generation. The results reveal that the shared bottom backbone helps visual stream learn fine-grained local features (#1 . #2), especially map information such as lane divider, pedestrian crossing, and lane boundary in our work. This is more effective for hard scenes, but to some extent, it affects the effectiveness of global features in simple scenes. For #3, the dual-input network involves a visual stream and a structural stream, which, beyond appearance information, can provide more detailed BEV structural information, thereby resulting in greater performance improvements in hard and semi-hard scenes. Based on the above discussion, a shared bottom backbone and a structural stream can maximize the robustness of global features under different environments.
BEV Types. We then perform ablation experiments to study the impact of varying types of BEV supervised models on VPR. Given the restricted variety of map elements within the nuScenes dataset, we focus on three key static elements: lane divider, pedestrian crossing, and lane boundary. As depicted in Table V, when training without structural information, the baseline achieves 88.93% R@1 and 4.58% R@1, respectively. When providing structural information of lane divider to the model, our method achieves gains of 2.32% and 9.75% on R@1 and R@1. More types of BEV consistently improves our method, particularly benefiting the categories of lane divider and lane boundary, which can be attributed to the higher proportion of road divider and road boundary in the image compared to pedestrian crossing. While three types of static elements with structural characteristics can achieve the best performance improvement.
@@ V CONCLUSIONS
In this paper, we propose a new BEV-enhanced VPR framework by exploiting the structural cues in BEV from a single monocular camera and introduce a new dataset for place recognition. The cores of our success lie in the introduction of BEV structural information and the promotion of the shared bottom backbone for local feature learning in the visual stream. Analytical experiments exhibit the superiority of BEV structural information in retrieving hard samples. We hope our work can enlighten more research on camera-based viusl place recognition.

### v2

@@ IV-A Dataset
The exclusive reliance on GPS for distinguishing between positive and negative samples presents challenges [30]. Images captured in close geographical location may fail to depict identical scenes when oriented differently, leading to pseudo positive samples. And variations in camera position while photographing the same site may introduce the possibility of pseudo negative samples. Due to the limitations in reflecting place position with camera coordinates, we employ image coordinates rather than camera coordinates to determine positive samples corresponding to a query sample.
In this work, the position of an image is defined as in front of the camera, aligned with the center of BEV image. After comparison, the visual effect of the positive images determined by this position is the best. In order to obtain more reasonable initial positive and negative samples, each image is represented by a binary tuple , where, represents the directional vector of camera-captured images, denoted as
where represents the position of the camera.
Then, we apply the Euclidean distance metric to calculate the distance between two images and those with a distance less than the threshold from other images are regarded as candidate positive samples, followed by angle calculation between the direction vectors of the query and positive samples to determine the final positive samples, i.e.,
where and represent the direction vectors of the query image and the candidate positive images respectively. To specify, the distance and angular thresholds for positive samples are set to and respectively.
Furthermore, in order to more clearly analyze the pros and cons of different modalities, we classify samples based on the appearance differences between query and positive images, and define it as follows (refer to Tab. I for details):
@@
• Hard Sample: significant illumination differences. For instance, the query image is taken during daylight, while all positive images are captured at nighttime.
@@
Partition of Training and Test Sets. Considering that NuScenes only involves four regions, each with data collected in relatively concentrated periods, we opt for a holistic partitioning approach for data across all areas. In VPR, the research on recalling from day to night, as well as from night to rain presents more challenges, but holds greater significance. Therefore, we aim to achieve a uniform distribution of such data in both the training and testing sets. This data distribution design facilitates a more comprehensive evaluation of the model’s adaptability to complex environmental changes, enhancing the model’s generalization performance.
We first establish an undirected graph with scenes as nodes and similarities between scenes as edge weights, where the scene similarity is defined as the proportion of each query image in one scene that has a corresponding positive image appearing in another scene. Specifically, if all query images have positive images in the same scene, the scene similarity is set to 1. After processing all the scenes, there are 32 isolated scenes without connecting to any other scenes, making them unusable for VPR but usable for BEV generation. Finally, we leverage the official API of NuScenes dataset to acquire scene labels, such as "night" or "day", and based on which, the data is manually balanced, generating the partitioning results in Tab. II. Note that there is no any instance and scene overlap between the training set and the test set. This dataset is called VPR-NuScenes.
@@ IV-B Implementation Details
Architecture. We implement BEV2PR in PyTorch framework and reproduce those methods with open-source codes for a fair comparison, including SPoC [25], NetVLAD [12], GeM [15], Conv-AP [18], EigenPlaces [26], MixVPR [27]. For all techniques involving BEV generation and VPR, we use EfficientNet-B0 as the backbone cropped at the last convolutional layer. And the sub-backbone in the visual stream is cropped from the block of EfficientNet-B0. Considering the information advantages of different scales, for BEV generation, we adopt a method of upsampling the output of the block and concatenating it with the output of the block as final extracted features. The Camera-to-BEV model outputs a map at a resolution with three classes, including lane boundary (Bound.), lane divider (Div.), and pedestrian crossing (P.C.), i.e., . Each input front-view image is resized from to in all experiments for all models.
Training. During Stage I, we train the Camera-to-BEV network with EfficientNet-B0 pre-trained on ImageNet for 40 epochs. The Adam optimizer has a weight decay of with a learning rate of . During Stage II, the final descriptor dimension of all methods is set to . We train the whole network using Adam optimizer having a weight decay of with a learning rate of for epochs. For each mini-batch, the number of positive samples , negative samples = 6 and are set for the triplet loss. In addition, the models for the whole dataset and all other subsets use the same hyper-parameters, with , , specifically set in the multi-head loss function.
@@
- | - | - | +1.60 | +0.63 | +0.69 | +0.91 | +0.13 | +0.02 | +10.03 | +13.18 | +16.05 | +3.14 | -0.63 | -0.42 |
’’ represents ’improvement’. Green numbers show our method’s performance improvement over the baseline, while blue numbers indicate the maximum performance improvement among various methods. Bold black numbers highlight the best performance of different methods. These notations are consistent in Tab. IV and Tab. V.
R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |||||
① | ✗ | ✗ | 88.93 | 93.87 | 94.80 | 94.48 | 98.62 | 99.16 | 4.58 | 8.88 | 12.03 | 84.83 | 94.98 | 97.49 | |||||
② | ✗ | ✓ | 88.75 | 93.73 | 94.90 | 93.70 | 97.79 | 98.50 | 10.89 | 19.48 | 26.93 | 86.51 | 95.61 | 97.28 | |||||
- | - | -0.18 | -0.14 | +0.10 | -0.78 | -0.83 | -0.66 | +6.31 | +10.60 | +14.90 | +1.68 | +0.63 | -0.21 | ||||||
③ | ✓ | ✗ | 90.59 | 94.03 | 95.45 | 95.24 | 98.21 | 99.14 | 15.19 | 22.64 | 29.23 | 89.12 | 94.04 | 96.55 | |||||
- | - | +1.66 | +0.16 | +0.65 | +0.76 | -0.41 | -0.02 | +10.61 | +13.76 | +17.20 | +4.29 | -0.94 | -0.94 | ||||||
④ | ✓ | ✓ | 91.40 | 95.53 | 96.85 | 95.67 | 98.46 | 99.03 | 22.64 | 42.41 | 54.44 | 89.85 | 96.65 | 98.74 | |||||
- | - | +2.47 | +1.66 | +2.05 | +1.19 | -0.16 | -0.13 | +18.06 | +33.53 | +42.41 | +5.02 | +1.67 | +1.25 |
① | - | - | - | 88.93 | 93.87 | 94.80 | 94.48 | 98.62 | 99.16 | 4.58 | 8.88 | 12.03 | 84.83 | 94.98 | 97.49 |
② | ✓ | - | - | 91.25 | 94.90 | 96.05 | 96.09 | 98.83 | 99.41 | 14.33 | 25.50 | 32.95 | 89.12 | 95.71 | 98.12 |
③ | - | ✓ | - | 89.96 | 94.21 | 95.56 | 95.19 | 98.36 | 99.08 | 9.46 | 22.35 | 30.66 | 86.72 | 94.56 | 97.28 |
④ | - | - | ✓ | 90.11 | 94.90 | 96.20 | 94.94 | 98.84 | 99.46 | 15.76 | 24.93 | 34.96 | 87.13 | 95.82 | 98.22 |
⑤ | ✓ | ✓ | - | 91.02 | 95.05 | 96.30 | 96.09 | 98.59 | 99.30 | 14.33 | 30.09 | 39.54 | 87.34 | 96.65 | 98.33 |
⑥ | ✓ | - | ✓ | 91.11 | 94.65 | 95.82 | 95.92 | 98.59 | 99.20 | 17.48 | 26.07 | 33.24 | 87.97 | 95.08 | 97.59 |
⑦ | - | ✓ | ✓ | 90.77 | 94.69 | 95.68 | 95.79 | 98.84 | 99.48 | 9.74 | 17.19 | 23.21 | 89.02 | 97.07 | 98.43 |
⑧ | ✓ | ✓ | ✓ | 91.40 | 95.53 | 96.85 | 95.67 | 98.46 | 99.03 | 22.64 | 42.41 | 54.44 | 89.85 | 96.65 | 98.74 |
Evaluation. The top- () recall is adopted as the evaluation metric, based on which, we evaluate our framework on the whole dataset as well as three subsets with different levels of recall difficulty, i.e., R@ (whole), R@ (easy), R@ (hard), R@ (semi-hard). To specify, we first calculate the similarity between sample and sample () in the test dataset/subset, based on which, the index of most similar samples is obtained, among which, we then exclude consecutive frames and detect positive samples, indicating successful recall.
@@ IV-C Evaluation for Visual Place Recognition
We apply six baseline models to our framework and evaluate the VPR performance based on monocular images. Tab. III shows the recall results on VPR-NuScenes. All methods incorporating structural streams have shown improvements in almost all metrics, with the absolute gains on R@1 over the whole dataset being 5.60%, 2.59%, 4.19%, 2.47%, 2.80%, 1.60%. From the results on other subsets, it can also be observed that datasets do not limit BEV2PR: on datasets with remarkable appearance differences, it can effectively improve performance (refer to next two paragraphs); on datasets with insignificant appearance differences, BEV2PR maintains original RGB performance. Note that we train all models incorporating structural streams with the same hyper-parameters (e.g., loss function coefficients). This may have prevented some models from achieving optimal performance, potentially explaining why MixVPR with the structural stream under-performs compared to ConvAP. However, this does not detract the overall performance improvements.
@@
Illumination Change. Compared to the easy set, the performance improvement on the hard set is more significant, with a maximum increase of 18.06% on R@1. This can be attributed to the fact that in scenes with significant illumination changes, the semantic and texture information becomes completely ineffective and structural features in the BEV frame are relatively stable. As shown in the second row of Fig. 3, it is difficult for human eyes to distinguish the complete scene, while just like in the nighttime scenes of autonomous driving in the real world, grounds are particularly clear compared to other areas, especially lane boundary and lane divider. Fig. 4 presents the feature maps of RGB and BEV during the day and night, respectively. It is evident that under such conditions, the effectiveness of structural information in the BEV frame is far greater than that of appearance information in the RGB frame.
@@
Image Clarity Change. The methods with added structural cues can also yield considerable performance enhancements in scenes where there is a substantial change in image clarity, with an average improvement of 5.65% on R@1 across different models. As shown in the third row of Fig. 3, compared with positive images, the query image is more blurry, but the overall image content remains basically unchanged. The semantic and texture information in certain areas is relatively complete and the structural information can further assist in generating more comprehensive features.
@@ IV-D Ablation Study
Shared Bottom Backbone. To demonstrate the effectiveness of the shared bottom backbone, we intentionally degenerate our framework into two versions, as depicted in Tab. IV, one (②) is single-stream with a bottom backbone pre-trained through BEV generation, the other (③) is dual-stream with a bottom backbone not pre-trained via BEV generation. The results reveal that the pre-trained bottom backbone helps visual stream learn fine-grained local features (① . ②), especially map information such as lane divider, pedestrian crossing, and lane boundary in our work. This is more effective for hard scenes, but to some extent, it affects the effectiveness of global features in simple scenes. For ③, the dual-stream network can provide more detailed BEV structural information, thereby resulting in greater performance improvements in hard and semi-hard scenes. Based on the above discussion, the dual-stream network with a shared bottom backbone can maximize the robustness of global features under different environments.
BEV Types. We then perform ablation experiments to study the impact of varying types of BEV supervised models on VPR. Given the restricted variety of map elements within the NuScenes dataset, we focus on three key static elements: lane divider, pedestrian crossing, and lane boundary. As depicted in Tab. V, when training without structural information, the baseline model achieves 88.93% R@1 and 4.58% R@1, respectively. When providing structural information of lane divider to the model, BEV2PR achieves gains of 2.32% and 9.75% on R@1 and R@1. More types of BEV consistently improves our method, particularly benefiting the categories of lane divider and lane boundary, which can be attributed to the higher proportion of road divider and road boundary in the image compared to pedestrian crossing. While three types of static elements with structural characteristics can achieve the best performance improvement.
Sequence | 00 | 02 | 05 | 06 | Mean |
Vision | 93.7 | 75.2 | 84.3 | 95.6 | 87.2 |
Vision + Structure | 94.4 | 78.1 | 85.9 | 98.2 | 89.2 |
IV-E Generalization.
We also evaluate the VPR performance of our methods on the KITTI dataset [31] utilizing models trained on VPR-NuScenes. The dataset partition of KITTI is the same as [4]. We implement BEV2PR with MixVPR as the aggregation module of the visual stream. Tab. VI presents the R@1 on the validation set. The method with added structural cues demonstrates more robust performance, although the BEV features may not be completely accurate.
@@ V CONCLUSIONS
In this paper, we propose a new BEV-enhanced VPR framework by exploiting the structural cues in BEV from a single monocular camera and introduce a new dataset for place recognition. The cores of our success lie in the introduction of BEV structural information and the promotion of the shared bottom backbone for local feature learning in the visual stream. Analytical experiments exhibit the superiority of BEV structural information in retrieving hard samples. We hope our work can enlighten more research on VPR. In future work, we will explore the impact of additional BEV categories on VPR performance and lightweight our dual-stream model through distillation techniques.
Acknowledgments
This work was supported in part by the Beijing Natural Science Foundation (Grant No. L223003, JQ22014), the Natural Science Foundation of China (Grant No. U22B2056, 62192782, 62036011, U2033210, 62102417), the Project of Beijing Science and technology Committee (Project No. Z231100005923046). Jin Gao was also supported in part by the Youth Innovation Promotion Association, CAS.